OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (anonymous):

Dan, I needed help on my last question still :(

OpenStudy (anonymous):

(6, 6)

OpenStudy (anonymous):

x+3/y^2

OpenStudy (anonymous):

9/36

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

oh 1/4

OpenStudy (anonymous):

But tahts not one of my options dan! :( on the last Q it wasnt one of the options either

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

y = +-4x + 12

OpenStudy (anonymous):

But dannnn ... :(

OpenStudy (anonymous):

y = (1/12)x^2 -12y = x^2 x = (1/12)y^2 y^2 = 6x

OpenStudy (anonymous):

what do i do to our equation that turns it into one of those?

OpenStudy (anonymous):

(3, -3)

OpenStudy (anonymous):

well it goes all the way to -3

OpenStudy (anonymous):

Darn it... :/ im sorry if i was confusing

OpenStudy (dan815):

start over, look up what a directrix means, give me a picture i can work with and ill tell you the formula

OpenStudy (anonymous):

okay one sec please dont go anywhere

OpenStudy (anonymous):

A parabola is defined as follows: For a given point, called the focus, and a given line not through the focus, called the directrix, a parabola is the locus of points such that the distance to the focus equals the distance to the directrix. @dan815

OpenStudy (dan815):

ok

OpenStudy (dan815):